OpenStudy (anonymous):

Please help! Thank You! Find the equation of the circle whose center and radius are given. center (0, 8), radius = 8 Equation answer should fill in these blanks __+(___)____

OpenStudy (anonymous):

\[(x-0)^2+(y-8)^2=8^2\\x^2+(y-8)^2=64 \]